There are 520 people in a village...200 of them are females what is the percentage of men?

Found 2 solutions by Boreal, stanbon:
Answer by Boreal(15235)

It is 320/520 or 61.5%
Alternatively, you can do 200/520 (females), get the percentage, and subtract it from 100%, since there are only two possibilities.

Answer by stanbon(75887)

There are 520 people in a village...200 of them are females what is the percentage of men?
====
Ans: (520-200)/520 = 320/520 = 32/52 = 16/26 = 8/13 = 0.62 = 62%
------------
Cheers,
Stan H.
-------------